Logiweb dictionary of append

Up Help

0 0 append
1 1 nani tetu nike...:1

2 2 append(∗, ∗)

The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-05-29.UTC:19:57:57.193456 = MJD-53519.TAI:19:58:29.193456 = LGT-4624113509193456e-6